# Conditionally valid Probabilistic Conformal Prediction

### v1

@@ 2 Conditionally valid Probabilistic Conformal Prediction
    We introduce $\lambda_{x,y}=\inf\left\{{t\in\mathbb{R}\colon y\in\mathcal{R}(x;t)}\right\}$, which can be considered as a conformity score for the input $x$ and the observation $y$. On the calibration dataset, we determine $\tau_{x}$ using the model $\Pi_{Y\mid X}$; see details later in (3). We also introduce an increasing function $f_{\tau}(\lambda)$ parameterized by $\tau>0$. An example of such a function is $f_{\tau}(\lambda)=\tau\lambda$. We then calculate the empirical measure $\mu$ of the transformed conformity values $f_{\tau_{X_{k}}}^{-1}(\lambda_{X_{k},Y_{k}})$. Finally, we determine the quantile $Q_{1-\alpha}(\mu)$ and construct the prediction set as follows
    $\mathcal{C}_{\alpha}(X_{n+1})=\mathcal{R}\left({X_{n+1};f_{\tau_{X_{n+1}}}\bigl{(}Q_{1-\alpha}(\mu)\bigr{)}}\right),$ (1)
    which achieves marginal validity $\mathbb{P}\left({Y_{n+1}\in\mathcal{C}_{\alpha}(X_{n+1})}\right)\geq 1-\alpha$, see Theorem 3.1.
Remark 2.1.
Conformal prediction methods are based on a conformity score $V(x,y)$, which evaluates how well the model prediction matches $y$. Given $V$, one can construct a family of confidence intervals $\mathcal{R}(x,t)=\{y\in\mathcal{Y}\colon V(x,y)\leq t\}$. With this definition, $\lambda_{x,y}=V(x,y)$; therefore, $\lambda_{x,y}$ plays the role of conformity measure. In (1), if we take $f_{\tau}(\lambda)=\lambda$, the standard conformal method is used. On the other hand, if we set $f_{\tau}(\lambda)=\tau$, then the confidence set is completely determined by the conditional generative model (we trust our conditional distribution but lose conformal guarantees). Taking $f_{\tau_{n+1}}(Q_{1-\alpha}(\mu))$ provides a compromise between the “simple” conformal method $Q_{1-\alpha}(\mu)$ and the confidence set derived from the conditional distribution, expressed by $\tau_{x}$. In this context, $\tilde{V}(x,y)=f_{\tau_{x}}^{-1}(\lambda_{x,y})$ can be viewed as an “adjusted” conformity measure. Consequently, (1) represents the conformal prediction set derived from the conformity score function $\tilde{V}$. This relationship directly establishes the finite sample marginal validity.
Remark 2.2.
If the predictive distribution has multiple modes, a single interval centered around the predictive mean often fails to provide an informative prediction set. Ideally, $\mathcal{R}(x;t)$ should correspond to the intervals with the highest probability density (HPD) of $\textup{P}_{Y\mid X}$; HPD regions are difficult to determine in practice, even when the conditional predictive density is available. [41] proposes a way to approximate HPD domains by using implicit conditional generative models. In this case, the prediction sets $\mathcal{R}_{z}(x;t)$ can depend on an exogenous variables $z\in\mathcal{Z}$, such as a union of balls with radius $t$ centered at points sampled from $\Pi_{Y\mid X}$.
2.1 The CP2 framework
We will now present our method for constructing adaptive prediction sets defined in (1). These sets are based on confidence set $\mathcal{R}_{z}(x;t)$ satisfying the following assumption.
H 1.
For all $(x,z)\in\mathbb{R}^{d}\times\mathcal{Z}$, the confidence sets $\{\mathcal{R}_{z}(x;t)\}_{t\in\mathbb{R}}$ are non-decreasing, and $\cap_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)=\emptyset$, $\cup_{t\in\mathbb{R}}\mathcal{R}_{z}(x;t)=\mathcal{Y}$, $\cap_{t^{\prime}>t}\mathcal{R}_{z}(x;t^{\prime})=\mathcal{R}_{z}(x;t)$.
In simpler terms, the size of $\mathcal{R}_{z}(x;t)$ grows with $t$. If we select a large enough value for $t$, we can cover the entire output space $\mathcal{Y}$. Let’s introduce the parameter $\lambda_{x,y,z}$ which corresponds to the minimal radius needed to guarantee that the confidence set contains the output value $y$
$\lambda_{x,y,z}=\inf\left\{{t\in\mathbb{R}\colon y\in\mathcal{R}_{z}(x;t)}\right\}.$ (2)
Lemma 2.3.
Assume 1 holds. For any $(x,y,z)\in\mathbb{R}^{d}\times\mathcal{Y}\times\mathcal{Z}$, $\lambda_{x,y,z}\in(-\infty,+\infty)$ and $y\in\mathcal{R}_{z}(x;\lambda_{x,y,z})$.
Our method relies on a family of transformations, denoted as $\{\lambda\mapsto f_{\tau}(\lambda)\}_{\tau\in\mathbb{R}}$, which balance the following two factors:
  • The optimal parameter $\lambda_{x,y,z}$ that ensures $y$ is included in the confidence set $\mathcal{R}_{z}(x;\lambda_{x,y,z})$.
  • The parameter $\tau_{x,z}$ obtained from the probabilistic model $\Pi_{Y\mid X=x}$.
H 2.
There exists $\varphi\in\mathbb{R}$ such that $\tau\in\mathbb{R}\mapsto f_{\tau}(\varphi)$ is increasing and bijective. Additionally, the function $\lambda\in\mathbb{R}\cup\{\infty\}\mapsto f_{\tau_{x,z}}(\lambda)$ is increasing for any $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$, where $\tau_{x,z}$ is defined in (3).
Given the significance level $\alpha\in[0,1]$, consider
$\tau_{x,z}=\inf\left\{{\tau\in\mathbb{R}\colon\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau}(\varphi)))\geq 1-\alpha}\right\}$ (3)
where by convention, we set $\inf\emptyset=-\infty$.
Lemma 2.4.
Assume 1-2 hold, and let $\alpha\in(0,1)$, $x\in\mathbb{R}^{d}$, $z\in\mathcal{Z}$. If $\Pi_{Y\mid X=x}$ is a probability measure, then $\tau_{x,z}\in(-\infty,\infty)$ and $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi)))\geq 1-\alpha$.
In words, $\tau_{x,z}$ ensures that the confidence set $\mathcal{R}_{z}(x;f_{\tau_{x,z}}(\varphi))$ is approximately conditionally valid when the distribution of $Y$ given $X=x$ is well approximated by the probabilistic model $\Pi_{Y\mid X=x}$.
For notational simplicity, set $\bar{\tau}_{k}:=\tau_{X_{k},Z_{k}}$ and $\bar{\lambda}_{k}:=\lambda_{X_{k},Y_{k},Z_{k}}$. Given $X_{n+1}\in\mathbb{R}^{d}$, we sample $Z_{n+1}\sim\Pi_{Z\mid X=X_{n+1}}$ and construct the resulting CP2 prediction set as
$\mathcal{C}_{\alpha}(X_{n+1})=\mathcal{R}_{Z_{n+1}}\left({X_{n+1};f_{\bar{\tau}_{n+1}}\bigl{(}Q_{1-\alpha}(\mu)\bigr{)}}\right),$ (4)
where $Q_{1-\alpha}(\mu)$ is the $1-\alpha$ quantile of the distribution $\mu$ is given by
$\mu=\frac{1}{n+1}\sum\nolimits_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$ (5)
Now, let’s explore two examples of the CP2 framework.
2.2 CP2 with Explicit Conditional Generative Model: CP2-HPD
We illustrate a specific instance of our framework, called CP2-HPD. We suggest using this approach when an approximate of the density function is known, denoted by $\gamma_{Y\mid X=x}$. The confidence set is defined as $\mathcal{R}(x;t)=\{y\in\mathcal{Y}\colon\gamma_{Y\mid X=x}(y)\geq t\}$. We omit the variable $z$ from the notation, as we do not consider exogenous randomization in this case. The parameter $\tau_{x}$ is obtained by solving $\Pi_{Y\mid X=x}(\mathcal{R}(x;f_{\tau}(\varphi)))=1-\alpha$. We then compute $\lambda_{x,y}=\gamma_{Y\mid X=x}(y)$ and derive the prediction set as
$\mathcal{C}_{\alpha}(x)=\left\{{y\in\mathcal{Y}\colon\gamma_{Y\mid X=x}(y)\geq f_{\tau_{x}}\left({Q_{1-\alpha}\left({\mu}\right)}\right)}\right\}.$
If we take $f_{\tau}(\lambda)=\lambda$ and $\varphi=1$, the method shares similarity with the CD-split method, proposed in [20]. While CD-split uses $\lambda_{x,y}=\gamma_{Y\mid X=x}(y)$ as the conformity score, our method uses $f_{\tau_{x}}^{-1}(\lambda_{x,y})$, which incorporates the information from $\tau_{x}$ to modify $\gamma_{Y\mid X=x}(y)$. Additionally, CP2-HPD does not rely on binning, unlike the CD-split approach.
2.3 CP2 with Implicit Conditional Generative Model: CP2-PCP
@@
We also develop a second instance of the CP2 algorithm, which is inspired by [41]. Unlike CP2-HPD, this approach does not require the conditional density. Instead, it is designed for cases where the conditional generative model (CGM) $\Pi_{Y\mid X}$ is implicit, meaning we cannot evaluate it pointwise while being able to sample from it. For each calibration point $X_{k}$, we draw $M$ random variables $\{\hat{Y}_{k,i}\}_{i=1}^{M}$ from $\Pi_{Y\mid X=X_{k}}$. We denote $Z_{k}=(\hat{Y}_{k,1},\cdots,\hat{Y}_{k,M})$ and consider the confidence sets as the union of spheres centered around the sample points:
$\mathcal{R}_{Z_{k}}(X_{k};t)=\cup_{i=1}^{M}\mathrm{B}(\hat{Y}_{k,i},t).$
With such choice, we get $\bar{\lambda}_{k}=\min_{i=1}^{M}\|Y_{k}-\hat{Y}_{k,i}\|$. We then draw a second sample $\{\tilde{Y}_{k,j}\}_{j=1}^{\tilde{M}}$, and compute $\bar{\tau}_{k}=\{\tau\in\mathbb{R}_{+}\colon\sum_{j=1}^{\tilde{M}}\mathds{1}_{\tilde{Y}_{k,j}\in\mathcal{R}_{Z_{k}}(X_{k};f_{\tau}(\varphi))}\geq 1-\alpha\}$. It can be verified that
@@
$\bar{\tau}_{k}=\left({\tau\mapsto f_{\tau}(\varphi)}\right)^{-1}\left\{{Q_{1-\alpha}\left({\frac{1}{\tilde{M}}\sum_{j=1}^{\tilde{M}}\delta_{\min_{i=1}^{M}\|\tilde{Y}_{j,k}-\hat{Y}_{k,i}\|}}\right)}\right\}.$
@@
$\mathcal{C}_{\alpha}(X_{n+1})=\left\{{y\in\mathcal{Y}\colon\min_{i=1}^{M}\|y-\hat{Y}_{n+1,i}\|\leq f_{\bar{\tau}_{n+1}}\left({Q_{1-\alpha}\left({\mu}\right)}\right)}\right\},$
@@
where $\mu$ is given in (5). The CP2-PCP method employs the same confidence set $\mathcal{R}_{z}(x;t)$ as the one used by PCP. This method effectively captures multimodalities using balls centered at likely outputs $\hat{Y}_{n+1,i}$. Furthermore, the conformity scores used by PCP correspond to our $\lambda_{x,y,z}$. However, the key distinction between the two algorithms lies in the additional parameter $\tau_{x,z}$ for CP2-PCP, which requires the generation of a second random sample from $\Pi_{Y\mid X=x}$.
Figure 3: Examples of $\bar{\tau}_{k}$ and $\bar{\lambda}_{k}$ for two different scenarios. In both cases, $\bar{\tau}_{k}$ is selected such that the probability of the corresponding confidence set $\mathcal{R}_{Z_{k}}(X_{k};f_{\bar{\tau}_{k}}(\varphi))$ under the distribution $\Pi_{Y\mid X=X_{k}}$ is equal to $1-\alpha$. On the other hand, $\bar{\lambda}_{k}$ is chosen as the smallest real number such that the true label $Y_{k}$ belongs to the confidence set $\mathcal{R}_{Z_{k}}(X_{k};\bar{\lambda}_{k})$.
2.4 Asymptotic Validity of CP2
To gain insights, we informally discuss the asymptotic conditional validity of CP2. We assume for simplicity that $\textup{P}_{Y\mid X}=\Pi_{Y\mid X}$, i.e., the predictive distribution is known. If 1 and 2 hold, then Lemma 2.3 shows that
$=\Pi_{Y\mid X=x}\left({\mathcal{R}_{z}(x;f_{\tau_{x,z}}(t))}\right).$
Note that $\Pi_{Y\mid X=x}(\mathcal{R}_{z}(x;f_{\tau_{x,z}}(t)))\geq 1-\alpha$ if and only if $t\geq\varphi$, which implies that
$\mathbb{P}(f_{\tau_{x,z}}^{-1}(\lambda_{x,Y,z})\leq t\mid X=x,Z=z)\geq 1-\alpha$   if and only if   $t\geq\varphi$. (6)
From (6) we deduce that the $(1-\alpha)$-quantile of $f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})$ is $\varphi$. The Glivenko–Cantelli Theorem [37, Theorem 19.1] demonstrates that $\sup_{t\in\mathbb{R}}|\mu(-\infty,t]-\mathbb{P}(f_{\tau_{X,Z}}^{-1}(\lambda_{X,Y,Z})\leq t)|\to 0$ almost surely as $n\to\infty$, where $\mu$ is defined in (5). Applying [37, Lemma 21.2], under weak conditions we deduce that $Q_{1-\alpha}(\mu)\to\varphi$ as $n\to\infty$.
If the quantile $Q_{1-\alpha}(\mu)$ is greater than $\varphi$, it indicates that there are insufficient calibration data points $Y_{k}\in\mathcal{R}_{Z_{k}}(X_{k};f_{\bar{\tau}_{k}}(\varphi))$. Conversely, if $Q_{1-\alpha}(\mu)$ is less than $\varphi$, it means that an excessive percentage of calibration data falls within the confidence set $\mathcal{R}_{Z_{k}}(X_{k};f_{\bar{\tau}_{k}}(\varphi))$. In this situation, the radius of the confidence set is increased to ensure the correct proportion of observations $Y_{k}$ covered.
Table 1: Confidence sets $\mathcal{R}(x;t)$ found in the literature and also discussed in [18, Table 1].
[24] [24] [22] $[\mathrm{pred}(x)-t,\mathrm{pred}(x)+t]$ $[\mathrm{pred}(x)-t\sigma(x),\mathrm{pred}(x)+t\sigma(x)]$ $(1+t)[q_{\alpha/2}(x),q_{1-\alpha/2}(x)]-tq_{1/2}(x)$ [9] [30] [34] $[q_{t}(x),q_{1-t}(x)]$ $[q_{\alpha/2}(x)-t,q_{1-\alpha/2}(x)+t]$ $[q_{\alpha/2}(x),q_{1-\alpha/2}(x)]\pm t(q_{1-\alpha/2}(x)-q_{\alpha/2}(x))$
2.5 Related Work
Approximating the conditional distribution of $Y\mid X$ to construct prediction sets has been extensively studied. Methods based on density estimation, such as [6, 25], achieve asymptotic validity under appropriate conditions. [19] uses kernel density estimation to construct asymmetric prediction bands. However, this method cannot handle bimodality as it generates a single interval. On the other hand, [35] partitions the domain of $Y$ into bins to create a histogram approximation of $\textup{P}_{Y\mid X}$. The authors showed that their method satisfies the marginal validity while achieving the asymptotic conditional coverage. This property has also been derived in the context of regression [24]. Asymptotic conditional coverage is also obtained in [34, 8] using quantile regression-based methods, or using cumulative distribution function estimators [20, 9]. Conditionally valid prediction sets have been shown to improve the robustness to perturbations [14]. [41] introduced a method adapted to distributions with separated high-density regions and with implicit CGM. Recently, [17] converted regression problems into classification problems by binning the range space and discretizing the labels. They learn an approximation of the conditional density to generate prediction sets that match the HPD regions. Similarly, [12] proposed a method that estimates the conditional density using neural network parameterized splines. The concept of nested sets was investigated in the work of [18]. Additionally, [32] proposed an approach leading to a confidence set $\mathcal{R}(x;\tau)=[\hat{q}_{\text{lo}}(x,\tau),\hat{q}_{\text{hi}}(x,\tau)]$, where $\hat{q}_{\text{lo}}(x,\tau)$ and $\hat{q}_{\text{hi}}(x,\tau)$ are quantile regressors for the conditional distribution $Y\mid X$. Moreover, $\tau$ is set such that at least $100(1-\alpha)\%$ of the calibration data fall within the prediction set.

### v2

@@ 2 The CP2 framework
    Finally, we need to introduce the conformity score for the considered setup. The natural choice is the minimal size of the set required to cover the observation $y$ at the input $x$ for the auxiliary variables $z$: $\lambda_{x,y,z}=\inf\left\{{t\in\mathsf{T}\colon y\in\mathcal{R}_{z}(x;t)}\right\}$.
  4. 4.
    The resulting procedure works as follows. For $k\in\{1,\dots,n\}$, we set $\bar{\tau}_{k}:=\tau_{X_{k},Z_{k}}$ and $\bar{\lambda}_{k}:=\lambda_{X_{k},Y_{k},Z_{k}}$ where $\{Z_{k}\}_{k=1}^{n}$ are sampled conditionally independently from $\bar{\Pi}_{Z\mid X=X_{k}}$. Given $X_{n+1}\in\mathbb{R}^{d}$, we sample $Z_{n+1}\sim\bar{\Pi}_{Z\mid X=X_{n+1}}$ conditionally independently from $\{(X_{k},Y_{k},Z_{k})\}_{k=1}^{n}$, and construct the resulting CP2 prediction set as
    $\mathcal{C}_{\alpha}(X_{n+1})=\mathcal{R}_{Z_{n+1}}\left({X_{n+1};f_{\bar{\tau}_{n+1}}\bigl{(}Q_{1-\alpha}(\mu_{n})\bigr{)}}\right),$ (4)
    where $Q_{1-\alpha}(\mu_{n})$ is the $1-\alpha$ quantile of the distribution $\mu_{n}$, and is given by
    $\mu_{n}=\frac{1}{n+1}\sum\nolimits_{k=1}^{n}\delta_{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})}+\frac{1}{n+1}\delta_{\infty}.$ (5)
    The transformation $\{v\mapsto f_{\tau}(v)\}_{\tau\in\mathsf{T}}$ balances the following two factors: (a) The optimal parameter $\lambda_{x,y,z}$ ensuring that $y$ is included in the confidence set $\mathcal{R}_{z}(x;\lambda_{x,y,z})$; (b) The parameter $\tau_{x,z}$ obtained from the probabilistic model $\Pi_{Y\mid X=x}$.
Table 1: Confidence sets $\mathcal{R}(x;t)$ found in the literature and also discussed in Gupta et al. (2022).
Lei et al. (2018) Lei et al. (2018) Kivaranovic et al. (2020) $[\mathrm{pred}(x)-t,\mathrm{pred}(x)+t]$ $[\mathrm{pred}(x)-t\sigma(x),\mathrm{pred}(x)+t\sigma(x)]$ $(1+t)[q_{\alpha/2}(x),q_{1-\alpha/2}(x)]-tq_{1/2}(x)$ Chernozhukov et al. (2021) Romano et al. (2019) Sesia & Candès (2020) $[q_{t}(x),q_{1-t}(x)]$ $[q_{\alpha/2}(x)-t,q_{1-\alpha/2}(x)+t]$ $[q_{\alpha/2}(x),q_{1-\alpha/2}(x)]\pm t(q_{1-\alpha/2}(x)-q_{\alpha/2}(x))$
We stress that CP2 is a general framework that can be adapted to many choices for conditional predictive density estimates, constructing the family of confidence sets, and selecting the calibration function $f_{\tau}(\lambda)$. However, we start with the simple example that shows that CP2 is more general than the classical split-conformal CP approach.
Figure 1: Predictions sets obtained via the standard CP and CP2 methods.
Simple example of CP2.
We begin with a simple application of CP2 to highlight its differences from the basic conformal approach, with $\mathcal{Y}=\mathbb{R}$. The calibration sets are defined as: $\mathcal{R}(x;t)=\{y\in\mathcal{Y}\colon|y-\text{pred}(x)|\leq t\}$ for $t\in\mathbb{R}_{+}$. There are no auxiliary variables $z$ in this case, so we omit $z$ from the notation. Assumption 1 is easily satisfied with $\mathsf{T}=\mathbb{R}_{+}$. We then take $f_{\tau}(\lambda)=\tau\lambda$, $\tau\in\mathbb{R}_{+}$ and $\varphi=1$: 2 is also satisfied. Note that $f^{-1}_{\tau}(\lambda)=\lambda/\tau$ for $\tau\in\mathbb{R}_{+}^{*}$. Using the CP2 approach, we find that $\lambda_{x,y}=|y-\mathrm{pred}(x)|$, which corresponds to a standard conformity score. The classical conformal prediction method defines the $1-\alpha$ quantile based on the associated empirical measure $\nu_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{\bar{\lambda}_{k}}+\frac{1}{n+1}\delta_{\infty}$, where $\bar{\lambda}_{k}=|Y_{k}-\mathrm{pred}(X_{k})|$. CP2 differs from the basic conformal approach by introducing $\tau_{x}=\operatornamewithlimits{\arg\min}\{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}([\mathrm{pred}(x)\pm\tau])\geq 1-\alpha\}$ as in (3), where $[\mathrm{pred}(x)\pm\tau]=[\mathrm{pred}(x)-\tau,\mathrm{pred}(x)+\tau]$. The prediction set becomes $[\mathrm{pred}(x)\pm f_{\tau_{x}}(Q_{1-\alpha}(\mu_{n}))]$, where $\mu_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{\bar{\lambda}_{k}/\tau_{k}}+\frac{1}{n+1}\delta_{\infty}$. In this setting, the choice of $\varphi$ is irrelevant. Take $\varphi>0$ and denote $\tau^{\varphi}_{x}=\operatornamewithlimits{\arg\min}\{\tau\in\mathsf{T}\colon\Pi_{Y\mid X=x}([\mathrm{pred}(x)\pm f_{\tau}(\varphi)])\geq 1-\alpha\}$. It is easily seen that $\tau_{x}^{\varphi}=\tau_{x}/\varphi$. The prediction set becomes $[\mathrm{pred}(x)\pm f_{\tau^{\varphi}_{x}}(Q_{1-\alpha}(\mu^{\varphi}_{n}))]$, where $\mu^{\varphi}_{n}=\frac{1}{n+1}\sum_{k=1}^{n}\delta_{\bar{\lambda}_{k}/\bar{\tau}_{k}^{\varphi}}+\frac{1}{n+1}\delta_{\infty}$ with $\bar{\tau}_{k}^{\varphi}=\tau^{\varphi}_{X_{k}}$. Note that $Q_{1-\alpha}(\mu^{\varphi}_{n}))=\varphi Q_{1-\alpha}(\mu_{n}))$ and thus $f_{\tau^{\varphi}_{x}}(Q_{1-\alpha}(\mu^{\varphi}_{n}))=f_{\tau_{x}}(Q_{1-\alpha}(\mu_{n}))$ showing that, the prediction set does not depend on the choice of $\varphi$. To illustrate the advantage of our method, in Figure 1 we present the prediction sets obtained with the classical CP method and CP2 in the case of a Neal’s funnel-shaped distribution in 2 dimensions; see Neal (2003, Section 9).
The natural approach for the general case is to use the conditional distribution $\Pi_{Y\mid X=x}$ or its estimate to find Highest Predictive Density (HPD) regions and calibrate their size with the help of CP2. We develop the respective general algorithm CP2-HPD in Appendix B.1. However, the procedures to find HPDs are usually highly non-trivial and we only investigate this approach experimentally for synthetic data; see Section 4.1. Next, we provide a specific implementation of our general CP2 framework that is universally applicable.
CP2 with Implicit Conditional Generative Model: CP2-PCP.
@@
We also develop a second instance of the CP2 algorithm, inspired by Wang et al. (2023). Unlike CP2-HPD, this approach does not require the conditional density. Instead, it is designed for cases where the conditional generative model (CGM) $\Pi_{Y\mid X}$ is implicit, meaning we cannot evaluate it pointwise while being able to sample from it. For each calibration point $X_{k}$, we draw $M$ random variables $\{\hat{Y}_{k,i}\}_{i=1}^{M}$ from $\Pi_{Y\mid X=X_{k}}$. We denote $Z_{k}=(\hat{Y}_{k,1},\ldots,\hat{Y}_{k,M})$ and consider the confidence sets as the union of spheres centered around the sample points $\mathcal{R}_{Z_{k}}(X_{k};t)=\cup_{i=1}^{M}\mathrm{B}(\hat{Y}_{k,i},t)$. With such choice, we get $\bar{\lambda}_{k}=\min_{i=1}^{M}\|Y_{k}-\hat{Y}_{k,i}\|$. We then draw a second sample $\{\tilde{Y}_{k,j}\}_{j=1}^{\tilde{M}}$, and compute $\bar{\tau}_{k}=\{t\in\mathbb{R}_{+}\colon\tilde{M}^{-1}\sum_{j=1}^{\tilde{M}}\mathds{1}_{\mathcal{R}_{Z_{k}}(X_{k};f_{t}(\varphi))}(\tilde{Y}_{k,j})\geq 1-\alpha\}$. It is easily seen that
@@
$\bar{\tau}_{k}=\left({t\mapsto f_{t}(\varphi)}\right)^{-1}\left\{{Q_{1-\alpha}\left({\frac{1}{\tilde{M}}\sum_{j=1}^{\tilde{M}}\delta_{\min_{i=1}^{M}\|\tilde{Y}_{k,j}-\hat{Y}_{k,i}\|}}\right)}\right\}.$
@@
$\mathcal{C}_{\alpha}(X_{n+1})=\left\{{y\in\mathcal{Y}\colon\min_{i=1}^{M}\|y-\hat{Y}_{n+1,i}\|\leq f_{\bar{\tau}_{n+1}}\left({Q_{1-\alpha}\left({\mu_{n}}\right)}\right)}\right\},$
@@
where $\mu_{n}$ is given in (5). The CP2-PCP method employs the same confidence set $\mathcal{R}_{z}(x;t)$ as the one used by PCP. This method effectively captures multimodalities using balls centered at likely outputs $\hat{Y}_{n+1,i}$. Furthermore, the conformity scores used by PCP correspond to our $\lambda_{x,y,z}$. However, the key distinction between the two algorithms lies in the additional parameter $\tau_{x,z}$ for CP2-PCP, which requires the generation of a second random sample from $\Pi_{Y\mid X=x}$. This method is especially useful when solving equation (3) is intractable. We summarize CP2-PCP in Algorithm 1.
Algorithm 1 CP2-PCP
Input: dataset $\{(X_{k},Y_{k})\}_{k\in[n]}$, significance level $\alpha$, conditional distribution $\Pi_{Y\mid X}$, function $f_{t}$.
// Compute the $(1-\alpha)$-quantile
for $k=1$ to $n$ do
     Sample $\{\hat{Y}_{k,i}\}_{i=1}^{M}$ and $\{\tilde{Y}_{k,j}\}_{j=1}^{\tilde{M}}$ from $\Pi_{Y\mid X=X_{k}}$
     Set $\bar{\lambda}_{k}=\min_{i=1}^{M}\|Y_{k}-\hat{Y}_{k,i}\|$
     Set $\bar{\tau}_{k}=(t\mapsto f_{t}(\varphi))^{-1}\{{\textstyle{Q_{1-\alpha}({\textstyle{\tilde{M}^{-1}\sum_{j=1}^{\tilde{M}}\delta_{\min_{i=1}^{M}\|\tilde{Y}_{k,j}-\hat{Y}_{k,i}\|}}})}}\}$
$Q_{1-\alpha}\left({\mu_{n}}\right)\leftarrow\lceil(1-\alpha)(n+1)\rceil$-th smallest value in $\{f_{\bar{\tau}_{k}}^{-1}(\bar{\lambda}_{k})\}_{k\in[n]}\cup\{\infty\}$
// Compute the prediction set for a new point $x\in\mathbb{R}^{d}$
Sample $z=\{\hat{Y}_{i}\}_{i=1}^{M}$ and $\{\tilde{Y}_{j}\}_{j=1}^{\tilde{M}}$ from $\Pi_{Y\mid X=x}$
Set $\tau_{x,z}=\left({t\mapsto f_{t}(\varphi)}\right)^{-1}\{{\textstyle{Q_{1-\alpha}({\textstyle{\tilde{M}^{-1}\sum_{j=1}^{\tilde{M}}\delta_{\min_{i=1}^{M}\|\tilde{Y}_{j}-\hat{Y}_{i}\|}}})}}\}$
Output: $\mathcal{C}_{\alpha}(x)=\cup_{i=1}^{M}\mathrm{B}(\hat{Y}_{i},f_{\tau_{x,z}}({\textstyle{Q_{1-\alpha}({\textstyle{\mu_{n}}})}}))$.